What is an expiation of a line that contains the point (-4,3) and is perpendicular to the graph of x-2y=8
Answer by stanbon(75887)

What is an equation of a line that contains the point (-4,3) and is perpendicular to the graph of x-2y=8
------
Find slope of the given equation::
2y = x-8
y = (1/2)x-4
slope = 1/2
-----
So, slope of a perpendicular line must be -2.
-----------------
Form of desirec equation: y = mx + b
Solve for "b":
3 = -2*-4 + b
3 = 8 + b
b = -5
---------------
Equation::
y = -2x - 5
=====================
Cheers,
Stan H.